ACTIVITY IN BASIC CALCULUS

BASIC RULES IN DERIVATIVE


  Complete the blanks of the given function below with a number (except 0 and 1) to create your own problem and find the derivative of the function. Show your complete solution to each problem.

 

1. f(x) = -4x5 + ______x-4- 2468




2. f (x) =____x-3- _____x1/4 - 12x


3.f(x)= ____ "\sqrt[4]{x}" 3 - "\xlongequal{}"


4.f (x) = "\xlongequal{}" - _____x2 + "\sqrt[4]{x}"


 


1
Expert's answer
2022-03-21T11:32:54-0400

1. "f(x)=-4x\u2075+2x^{-4}-2468"

"f'(x)=-20x\u2074-8x^{-5}"


2. "f(x)=3x^{-3}-4x^{\frac{1}{4}}-12x"

"f'(x)=-9x^{-4}-x^{\frac{-3}{4}}-12"


3. "f(x)=4x^{\frac{3}{4}}"

"f'(x)=3x^{\frac{-1}{4}}"


4. "f(x)=-2x²+x^{\frac{1}{4}}"

"f'(x)=-4x+\frac{1}{4}x^{\frac{-3}{4}}"